### v1

# White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency

Abstract
Social biases can manifest in language agency. For instance, White individuals and men are often described as “agentic” and achievement-oriented, whereas Black individuals and women are frequently described as “communal” and as assisting roles. This study establishes agency as an important aspect of studying social biases in both human-written and Large Language Model (LLM)-generated texts. To accurately measure “language agency” at sentence level, we propose a Language Agency Classification dataset to train reliable agency classifiers. We then use an agency classifier to reveal notable language agency biases in $6$ datasets of human- or LLM-written texts, including biographies, professor reviews, and reference letters. While most prior NLP research on agency biases focused on single dimensions, we comprehensively explore language agency biases in gender, race, and intersectional identities. We observe that (1) language agency biases in human-written texts align with real-world social observations; (2) LLM-generated texts demonstrate remarkably higher levels of language agency bias than human-written texts; and (3) critical biases in language agency target people of minority groups—for instance, languages used to describe Black females exhibit the lowest level of agency across datasets. Our findings reveal intricate social biases in human- and LLM-written texts through the lens of language agency, warning against using LLM generations in social contexts without scrutiny.
White Men Lead, Black Women Help: Uncovering Gender, Racial, and Intersectional Bias in Language Agency
Yixin Wan and Kai-Wei Chang {elaine1wan, kwchang}@cs.ucla.edu
1 Introduction
Social biases manifest through different levels of agency in texts describing individuals of different demographic groups, such as race and gender Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). This is often through portraying minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020).
To measure agency level in texts, previous works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022) mostly utilized string matching for words in agentic and communal lexicons. However, existing tools like string matching and sentiment classifiers often fail to capture language agency. Table 11 in Appendix B shows that string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively. Table 1 provides a qualitative example: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words for both texts; sentiment classifier labels both sentences as “positive”, failing to identify implicit bias in agency.  Wan et al. (2023) utilized a model-based method for measuring agency and related gender biases, but only in the context of reference letters. Most aforementioned studies also only focused on single dimensions of biases and single types of text, limiting the scope of analysis.
Occupation Biography String Match Sentiment Classifier Agency Classifier
# Agen. # Comm.
Male Software Engineer Growing up in Vienna, he previously co-founded a legal marketplace startup connecting attorneys with potential clients, and has accumulated over 10 years of experience in iOS security research and app development, leading development on several high-profile projects such as … 0 0 Positive Agentic
Female Software Engineer She is passionate about helping developers build engaging, useful experiences and services that make people‚ lives better. 0 0 Positive Communal
Table 1: Example of gender biases manifested in language agency. Agentic languages are highlighted in blue, and communal languages are highlighted in red. String matching method (columns 3 and 4) and sentiment-based approach (column 5) fail to capture the bias in language style, whereas our proposed language agency classifier successfully identifies gender bias in agency.
Our study refines and extends the model-based approach to capture language agency biases in a broader scope of human-written and LLM-generated texts. In addition to studying both gender and racial biases, we propose to investigate intersectional biases in language agency. We first contribute a sentence-level Language Agency Classification dataset for training effective agency classifiers. Utilizing an agency classifier trained on our dataset, we reveal remarkable gender, racial, and intersectional biases in language agency on $6$ human-written and LLM-synthesized datasets, including biographies, professor reviews, and reference letters.
Results show that: (1) our findings on biases in human-written texts agree with previous social science discoveries. For instance, biographies of individuals with male-dominant occupations—such as pastor, architect, and software engineer—demonstrate greater extents of gender bias. This proves language agency as a reliable tool for studying biases in texts. (2) LLM-generated texts demonstrate notably greater levels of language agency gender bias, compared to human-written texts. (3) Both human-written and LLM-generated texts demonstrate remarkably higher language agency when describing dominant social groups—such as White men—compared to texts about minority groups—such as Black women, revealing severe biases in language agency. For example, reviews for Black professors—especially Black female professors—have the lowest language agency levels among faculties of all races. Black individuals also receive recommendation letters with the lowest language agency level, while White people—especially White males—receive the most agentic letters of recommendation. This study unveils critical language agency biases in human- and LLM-written biographies, professor reviews, and reference letters. Our findings warn against using model-generated texts while overlooking potential risks and harms in society.  [^1] Dataset and model checkpoints will be released upon acceptance.
2 Related Work
2.1 Biases in Human-Written Texts
The widespread presence of gender, racial, and intersectional bias in human society has significantly impacted human language Blodgett et al. (2020); Doughman et al. (2021). Since generative LLMs use extensive texts for training, understanding biases in texts is crucial for identifying and preventing biases in language technologies.
Bias in BiographiesWagner et al. (2016); Field et al. (2022), and  Park et al. (2021) studied gender biases in Wikipedia biographies.  Park et al. (2021) analyzed biases in power, agency, and sentiment words in biography pages;  Wagner et al. (2016) revealed negative linguistic biases in womens’ pages.  Field et al. (2022) and  Adams et al. (2019) studied racial biases in editorial traits such as length and academic rank.  Field et al. (2022); Adams et al. (2019) and  Lemieux et al. (2023) stressed the importance of studying intersectional gender and racial biases in Wikipedia. Along similar lines, Otterbacher (2015) found that IMDB biographies for Black female actresses describe them with more heavily subjective adjectives.
Bias in Professor ReviewsSchmidt visualized the extent of gendered language in RateMyProfessor reviews by string matching for gender-indicative words. Their observations align with prior findings on gender biases in student ratings for professors, revealing that instructors with female perceived gender received lower ratings than males (Roper, 2019; Macnell et al., 2014).  Reid (2010) showed that professors from racial minority groups—specifically Black and Asian groups—received more negative RateMyProfessor evaluations on overall quality than White faculties.  Chávez and Mitchell (2020) explored intersectional gender and racial biases in professor reviews, and found that female professors from racial minority groups received lower student evaluation scores compared to white males.
Bias in Reference Letters    Trix and Psenka (2003); Cugno (2020); Madera et al. (2009); Khan et al. (2021); Liu et al. (2009); Madera et al. (2019), and Wan et al. (2023) uncovered gender biases in letters of recommendation. For instance,  Trix and Psenka (2003); Madera et al. (2009) and  Madera et al. (2019) studied bias in the “exellency” of language.  Morgan et al. (2013); Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Polanco-Santana et al. (2021); Chapman et al. (2022); Girgis et al. (2023) investigated racial biases in reference letters:  Girgis et al. (2023) studied biases in emotional words and language traits like tone, but did not open-source their evaluation tools;  Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Chapman et al. (2022); Polanco-Santana et al. (2021), and Chapman et al. (2022) used string matching for word-level bias analysis. For example,  Powers et al. (2020) and  Chapman et al. (2022) showed that racial minority groups are significantly less frequently described with standout words than their White colleagues.
Most above-mentioned works, however, studied biases in simple language traits like length, words, or sentiments (e.g. excellency, tone), which often fail to capture biases in intricate language styles.
2.2 Bias in Language Agency
An increasing body of recent studies have investigated biases in intricate language styles, such as language agency Sap et al. (2017); Ma et al. (2020); Stahl et al. (2022); Wan et al. (2023).
Agentic language describes “proactive” characteristics such as speaking assertively, influencing others, and initiating tasks. Communal language portrays “reactive” characteristics like caring for others, providing assistance, and sustaining relationships (Madera et al., 2009; Wan et al., 2023). Bias in language agency can arise due to the disproportionate portrayal of agency and power, e.g. by showing women as submissive and powerless and men as assertive and dominant (Stahl et al., 2022), or by describing racial minority groups with more communal language than agentic (Grimm et al., 2020; Polanco-Santana et al., 2021).
Evaluation    Akos and Kretchmar (2016); Sap et al. (2017); Ma et al. (2020); Grimm et al. (2020); Polanco-Santana et al. (2021); Park et al. (2021), and Stahl et al. (2022) measured language agency by string matching for agentic and communal verbs, and then calculate their occurrence frequencies. However, string-matching methods fail to consider the diversity and complexity of language, and could not capture implicit indicators of language agency, as illustrated in Figure 1.  Wan et al. (2023) was the first to adopt a model-based method to measure language agency gender biases in LLM-generated reference letters. Nevertheless, the scope of their analysis is constrained to the LLM-synthesized letters of recommendation.
We believe that language agency is a common theme over a broad scope of social biases across all types of human-written and model-generated texts. Therefore, we propose to explore gender, racial, and intersectional biases in language agency in a broader context.
3 Language Agency Bias
3.1 Definition
We define “language agency bias” to be the unequal representation of language agency in texts depicting people of different demographic identities, such as race and gender. Given a piece of natural language text, we calculate the agency of each sentence, and then measure language agency level as the percentage of agentic sentences within the text. Biases can then be manifested as harmful differences between the agency levels of texts about people of different genders or racial groups. Different from previous works (Sap et al., 2017; Ma et al., 2020; Stahl et al., 2022), which used string matching for agentic and communal words from pre-defined lexicons, we adopt a model-based approach through a language agency classification model, which can more accurately capture implicit language agency traits in texts. Different from Wan et al. (2023), we propose a sentence-level language agency classification dataset, and extend the scope of analysis for language agency bias to a broader scope of natural language texts, including personal biographies, professor reviews, and reference letters.
3.2 Language Agency Classification Dataset
We propose the Language Agency Classification (LAC) dataset, a ChatGPT (OpenAI, 2022)-generated and human-reannotated dataset, to build reliable and accurate agency classifiers for identifying biases in language agency.
3.2.1 Dataset Collection
To collect agentic and communal sentences, we first preprocess a personal biography dataset (Lebret et al., 2016) into sentences, then use ChatGPT to paraphrase each sentence into an agentic version and a communal version. Next, we invite $2$ human annotators to re-label each data as “agentic”, “communal”, or “neutral”. Full details for constructing the LAC dataset are in Appendix A.
3.2.2 Dataset Statistics
Inter-Annotator Agreement   We consider ChatGPT’s categorization of agentic and communal paraphrases as its classification labels. Then, we calculate Fleiss’s Kappa score Feinstein and Cicchetti (1990) between ChatGPT’s labeling and the two main human annotators. The reported Fleiss’s Kappa score is $0.71$, proving the satisfactory quality of our proposed LAC dataset. Observing a significantly smaller number of “neutral” entries compared to the $2$ other labels, we remove all “neutral” entries in the dataset, achieving a new Fleiss’s Kappa score of $0.90$ for the final LAC dataset.
Dataset Split   We split the annotated and aggregated dataset into Train, Test, and Validation sets with a $0.8$, $0.1$, $0.1$ ratio. Detailed statistics of each split are in Appendix A, Table 9.
4 Experiments
4.1 Training Language Agency Classifiers
We experiment with both discriminative and generative models as base models for training language agency classifiers. For discriminative model, we experiment with BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019). For generative model, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. (2023). Appendix B provides details of training and inferencing the classifiers, in which Table 11 reports classifier performances. Based on model performance on LAC dataset’s test set, we choose the fine-tuned BERT model as the language agency classifier in further experiments.
4.2 Revealing Language Agency Biases
4.2.1 Datasets
We experimented with $6$ datasets—including biographies, professor reviews, and reference letters—to reveal gender, racial, and intersectional biases through the lens of language agency. $3$ of the datasets were publicly accessible; due to a lack of data with racial information, we use ChatGPT to synthetically generate the rest $3$ of the datasets with racial information. Details of all datasets are in Appendix C, Table 12.
Personal Biographies   To study language agency gender biases, we use Bias in Bios De-Arteaga et al. (2019), a biography dataset extracted from Wikipedia pages. Since the biography data for different professions are significantly imbalanced, we randomly sample $120$ biographies for each gender for each of the professions. A full list of professions in the pre-processed dataset is in Appendix C, Table 13. Since the dataset of Wikipedia-based human-written biographies does not include racial information, we follow the Context-Less Generation (CLG) setting in  Wan et al. (2023) and utilize ChatGPT to generate a biographies dataset for males and females of $4$ racial groups: White, Black, Hispanic, and Asian. We then use this dataset to study gender, racial, and intersectional language agency biases in LLM-generated biographies.
Professor Reviews   We investigate language agency bias in human-written professor reviews using an open-access sample dataset of student-written reviews for professors  [^2] https://github.com/x-zhe/RateMyProfessor_Sample_Dataset, which was web-crawled from the RateMyProfessor website  [^3] https://www.ratemyprofessors.com/. We begin preprocessing by removing the majority of data entries without professors’ gender information. Since the remaining data is scarce and unevenly distributed across genders and departments, we remove data from departments with less than $10$ reviews for male or female professors. No further re-sampling or balancing is conducted on the remaining dataset, since that would remove most data entries. A full list of departments and corresponding gender distributions of professor reviews in the pre-processed dataset is provided in Appendix C, Table 14. Since the RateMyProfessor dataset does not include racial information, we again used ChatGPT under the CLG setting to generate a professor review dataset with synthetic comments for male and female professors of the $4$ above-mentioned racial groups, and from the same set of departments as in the RateMyProfessor data. Then, we use the dataset to comprehensively study language agency biases in model-synthesized professor reviews.
Reference Letters   We analyze language agency gender bias using the ChatGPT-generated reference letter dataset in  Wan et al. (2023)’s work. Since  Wan et al. (2023)’s dataset does not include racial information, we again adopted the CLG approach and employed ChatGPT to generate a reference letter dataset for individuals, with both gender and racial specifications. We then used this synthetic dataset to investigate language agency bias in gender, race, and intersectional identities in LLM-generated letters of recommendation.
4.2.2 Reported Metrics
For each dataset for experiments, we first calculate the average percentage of agentic and communal sentences in texts describing different social groups. We then mainly report the average ratio gap between agentic and communal sentences, to better reflect the overall language agency level. For gender bias evaluation, we further conducted statistical t-testing with alternative hypothesis ($H_{1}$) being that male texts are higher in agency level than female texts, and null hypothesis ($H_{0}$) being no significant agency differences. Statistical testing results are in Appendix D, Table 18.
4.3 Results
Below, we illustrate experiment results revealing different aspects of language agency biases. All results are averaged on random seeds 00, $1$, and $2$.
4.3.1 Gender Bias
Table 2 shows results of gender biason the $6$ human-written and LLM-generated datasets. Across all datasets, texts describing males are remarkably higher in language agency level than those describing females. Additionally, language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that language agency is effective in capturing biases in implicit language styles. Furthermore, both ratio-based (Table 2) and statistical testing results (Appendix D Table 18) show that LLM-generated texts demonstrate more severe language agency gender biases than human-written texts. To better visualize overall results on the $3$ LLM-generated datasets, we use $3$ Kernel Density Estimate graphs in the upper row of Figure 2. We observe that texts about females possess higher density to the lower language agency end (i.e. left), whereas texts about males have higher density to the higher agency level end (i.e. right).
Figure 1: Visualization of the average ratio gap between agentic and communal sentences for different genders, in the 8 stratification aspects of Bias in Bios and RateMyProfessor with most significant gender biases.
Dataset Gender AvgAvgAvg
Agen. Comm. Gap
Biography (Human) Female $63.81$ $36.19$ $27.61$
Male 68.87 31.13 37.73
Biography (LLM) Female $38.28$ $61.72$ $-23.45$
Male 42.52 57.48 -14.96
Professor Review (Human) Female $77.84$ $22.16$ $55.67$
Male 78.76 21.24 57.53
Professor Review (LLM) Female $32.79$ $67.21$ $-34.42$
Male 36.07 63.93 -27.85
Reference Letter (Wan et al.) Female $55.15$ $44.85$ $10.30$
Male 57.47 42.53 14.94
Reference Letter (LLM) Female $53.26$ $46.74$ $6.52$
Male 57.92 42.08 15.85
Table 2: Experiment results for language agency gender bias in natural language texts. Results with the highest language agency level for each dataset are in bold. Overall, texts describing males demonstrate a higher level of language agency than texts describing females.
Bias in Biography   We investigate gender biases in both human-written and LLM-generated biographies, and observe that the average ratio gap between agentic and communal sentences in male biographies is higher than that of females. Our observations align with findings in prior social science work (Wagner et al., 2016), which revealed gender bias towards females in Wikipedia biographies. We further stratify analysis on the human-written Bias in Bios dataset based on the professions of individuals in biographies in Appendix D Table 20, and visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 1. Drastic language agency gender bias are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Observations on language agency gender bias in biographies align with social biases, showing the effectiveness of the proposed language agency approach in capturing biases in implicit language traits.
Bias in Professor Review   We study gender biases in both human-written and LLM-generated reviews for professors. On both human- and LLM-written professor reviews, we discover that reviews for male faculties possess higher agency than those for female faculties, aligning with gender bias observations on real-world professor ratings (Schmidt, ; Roper, 2019; Macnell et al., 2014). We further stratify our analysis on the human-written RateMyProfessor dataset based on the professors’ departments in Appendix D Table 19, and visualize the top $8$ most biased departments in Figure 1. The greatest language agency biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated (200, 2009; Girgus, ; Seijo, ). Again, gender bias in language agency observed in our study aligns with real-world biases, further validating the effectiveness of language agency in capturing biases.
Bias in Reference Letter   On both Wan et al. (2023)’s dataset and our LLM-generated reference letter dataset, we observe that the average ratio gap between agentic and communal sentences in letters for male candidates is higher than that for female candidates. This observation aligns with prior works’(Madera et al., 2009, 2019; Khan et al., 2021; Wan et al., 2023) conclusions, pointing towards significant language agency gender biases towards females in letters of recommendation.
Dataset Race AvgAvgAvg
Agen. Comm. Gap
Biography (LLM) White $41.81$ $58.19$ $-16.38$
Black $36.39$ $63.61$ $-27.22$
Hispanic $39.06$ $60.94$ $-21.88$
Asian 44.33 55.67 -11.34
Professor Review (LLM) White $34.62$ $65.38$ $-30.76$
Black $31.35$ $68.65$ $-37.30$
Hispanic $35.84$ $64.16$ $-28.33$
Asian 35.92 64.08 -28.16
Reference Letter (LLM) White 57.50 42.50 15.00
Black $54.54$ $45.46$ $9.08$
Hispanic $55.31$ $44.69$ $10.63$
Asian $55.01$ $44.99$ $10.03$
Table 3: Experiment results for language agency racial bias in natural language texts. Highest language agency level for each dataset is in bold. Texts describing Black individuals have the lowest language agency level compared with other racial groups.
Figure 2: Kernel Density Estimate (KDE) plot of the average ratio gap between agentic and communal sentences on the 6 investigated datasets. The 3 figures in the upper row illustrate language agency gender bias, in which texts describing males have greater density towards the higher language agency end (i.e. to the right). The 3 figures in the lower row illustrate language agency racial bias, in which texts describing Black individuals tend to have greater density towards the lower language agency end (i.e. to the left).
4.3.2 Racial Bias
Table  3 shows experiment results for racial biases in language agency across $3$ LLM-generated datasets with racial information in our study. Overall, we observe that (1) reference letters possess a larger proportion of agentic sentences, whereas biographies and professor reviews mostly contain more communal descriptions, and (2) on all investigated datasets, language describing Black individuals has the lowest level of agency than that for other racial groups. We visualize experiment results on each dataset using $3$ Kernel Density Estimate graphs in the lower row of Figure 2. Across datasets, texts about Black individuals have a higher density than other racial groups in the lower language agency end (i.e. left), aligning with observations in quantitative results.
Bias in Biography   On the LLM-generated biography dataset, biographies of Asian and White individuals demonstrate the highest level of language agency. Shockingly, there is a $15.88$ difference between the agentic-communal ratio gaps for Asian biographies and Black biographies, indicating a strong racial bias towards Black individuals.
Bias in Professor Review   Results on the LLM-synthesized professor review dataset also show biases towards Black faculties. A $9.14$ discrepancy is observed between agentic-communal ratio gaps of reviews for Asian professors and for Black professors. Interestingly, studies on real-world professor ratings also found that Black professors received more negative reviews from students (Reid, 2010).
Bias in Reference Letter   We observe that highest language agency levels are observed in LLM-generated reference letters for White individuals, whereas Black individuals receive letters with the lowest language agency level. Agentic-communal ratio gap of reference letters for White individuals is $5.92$ higher than Black individuals, $4.37$ higher than Hispanic individuals, and $4.97$ higher than Asian individuals. Such racial biases in LLM-written letters of recommendation also aligns with previous social science findings (Powers et al., 2020; Chapman et al., 2022).
4.3.3 Intersectional Bias
Different from prior works, which mostly focused on single bias aspects, we are amongst the first to investigate intersectional biases in LLM-generated texts through the lens of language agency. Specifically, we use the $3$ LLM-generated datasets—which contains both gender and racial information—to study intersectional gender and racial biases. Results show that texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. To visualize results, we plot the average ratio gap between agentic and communal sentences in texts describing different genders for different racial groups as overlap horizontal bar graphs in Figure 1. Both quantitative results and visualized illustrations show intersectional racial and gender biases across 3 datasets.
Figure 3: Visualization of the average ratio gap between agentic and communal sentences in the 3 datasets for different intersectional genders and racial groups. Agency levels of texts describing males are notably higher than those describing females in each racial group.
Dataset Race Gender AvgAvgAvg
Agen. Comm. Gap
Biographies (LLM) White Male $43.87$ $56.13$ $-12.27$
Female $39.76$ $60.24$ $-20.49$
Black Male $37.23$ $62.77$ $-25.53$
Female $35.55$ $64.45$ $-28.91$
Hispanic Male $42.30$ $57.70$ $-15.40$
Female $35.82$ $64.18$ $-28.36$
Asian Male 46.68 53.32 -6.64
Female $41.98$ $58.02$ $-16.03$
Table 4: Experiment results for intersectional gender and racial bias in language agency in biographies. Highest language agency level for each dataset is in bold. Biographies for Black females are lowest in language agency level, whereas biographies for Asian males are highest in agency.
Bias in Biographies   Table 4 shows results for intersectional biases in LLM-generated biographies. We observe that (1) biographies of Asian males demonstrate the highest level of language agency, (2) biographies of Black females show the lowest level of agency, and (3) within each racial group, language agency is notably higher in male biographies than in female ones. Similar findings have also been made in social science works, unveiling intersectional racial and gender biases in biographies (Field et al., 2022; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015)
Professor Review   Results for intersectional biases in LLM-synthesized professor reviews are shown in Table 5 While reviews for Hispanic male professors have the highest language agency, reviews for Black female professors again show the lowest agency level. This aligns with findings of prior social science studies, which showed that female professors from racial minority groups tend to receive lower scores in student ratings (Chávez and Mitchell, 2020). Additionally, within racial groups, reviews for female professors have lower levels of language agency than male faculties.
Dataset Race Gender AvgAvgAvg
Agen. Comm. Gap
Professor Review (LLM) White Male $36.01$ $63.99$ $-27.98$
Female $33.23$ $66.77$ $-33.54$
Black Male $31.95$ $68.05$ $-36.10$
Female $30.74$ $69.26$ $-38.51$
Hispanic Male 38.52 61.48 -22.95
Female $33.15$ $66.85$ $-33.71$
Asian Male $37.81$ $62.19$ $-24.39$
Female $34.03$ $65.97$ $-31.93$
Table 5: Experiment results for intersectional gender and racial bias in language agency in professor reviews. Highest language agency level for each dataset is in bold. Reviews for Black female professors demonstrate the lowest language agency, while reviews for Hispanic male faculties have the highest agency level.
Dataset Race Gender Avg % Avg % Avg
Agen. Comm. Gap
Reference Letter (LLM) White Male 59.88 40.12 19.75
Female $55.12$ $44.88$ $10.24$
Black Male $56.74$ $43.26$ $13.49$
Female $52.34$ $47.66$ $4.67$
Hispanic Male $57.64$ $42.36$ $15.27$
Female $52.99$ $47.01$ $5.98$
Asian Male $57.44$ $42.56$ $14.87$
Female $52.59$ $47.41$ $5.19$
Table 6: Experiment results for intersectional gender and racial bias in language agency in reference letters. Highest language agency level for each dataset is in bold. Recommendation letters for Black females again are the lowest in language agency; letters for White males are the highest in agency level.
Reference Letters   Table 6 shows results for intersectional biases in LLM-generated reference letters. Results show that (1) reference letters for white males have the highest levels of language agency, (2) letters for Black females again demonstrate the lowest level of agency, and (3) within the same racial group, level of language agency is remarkably lower in letters of recommendation for females than those for males.
5 Conclusion
In this work, we propose and study gender, racial, and intersectional biases in language agency in a wide scope of human-written and LLM-generated texts. First, we contribute the Language Agency Classification (LAC) dataset for training accurate classifiers to measure language agency. We then utilize an agency classifier to reveal critical language agency biases in $6$ human-written and LLM-generated datasets, including biographies, professor reviews, and letters of recommendation. Our findings show that (1) language agency is effective in capturing gender, racial, and intersectional biases in texts, (2) while both are biased, LLM-generated texts demonstrate notably higher levels of language agency bias than human-written texts, and (3) severe biases in language agency target females and Black individuals, especially for Black females. Our findings point towards the imminence to further explore biases in language agency, and forewarn potential social risks of using LLM-generated texts without scrutiny.
Limitations
We identify some limitations of our study. First, due to the limited information within the datasets available for our study, we were only able to consider the binary gender and $4$ racial groups for bias analyses. However, we note that it is important and significant for further works to extend the investigation of the fairness problem in our study to other gender and racial minority groups. Second, our study only includes experiments on investigating language agency biases in personal biographies, professor reviews, and reference letters, among which $4$ were synthesized and only $2$ contained human-written texts. We encourage future studies to extend the exploration of language agency biases in broader domains of human-written texts. Lastly, due to cost and resource constraints, we were not able to further extend our experiments to larger synthesized datasets. Future works should be devoted to comprehensively evaluating biases from various data sources.
Ethics Statement
Experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
Appendix A Language Agency Classification (LAC) Dataset Construction
A.1 Preprocessing
For the base dataset, we utilize the “evaluation” split of WikiBio Lebret et al. (2016), a personal biography dataset with information extracted from Wikipedia. We preprocess the dataset by splitting each personal biography into sentences. To ensure that each sentence is informative and depicts the owner of the biography, we remove the first two sentences and the last sentence, which usually provide the birth date and the current status of the owners without describing their characteristics.
A.2 ChatGPT Generation
For each of the pre-processed sentences in personal biographies, we prompt the gpt-3.5-turbo-1106 version of ChatGPT with one-shot example Wang et al. (2020) to paraphrase it into an agentic version and a communal version. Specific prompt used in the dataset generation process is provided in Table 7. This guarantees the balance of the constructed dataset and prevents the propagation of pre-existing biases in the classifier training process.
Prompt You will rephrase a sentence two times to demonstrate agentic and communal language traits respectively. ’agentic’ is defined as more achievement-oriented, and ’communal’ is defined as more social or service-oriented. Example of agentic description: {}. Example of communal description: {}. Output your answer in a json format with two keys, ’agentic’ and ’communal’. The sentence is: ’{}’
Agentic Example [Name] is an achievement-oriented individual with 7 years of experience being in charge of people and projects in previous workplace environments.
Communal Example [Name] is a people- oriented individual with 7 years of experience being a part of various financial teams and projects in previous workplace environments.
Table 7: Prompt for synthesizing the Language Agency Classification dataset using ChatGPT.
A.3 Human Re-Annotation
In order to ensure the quality of data generation by ChatGPT, we invite two expert human annotators to label the generated dataset. Both human annotators are native English speakers, and volunteered to participate in this study. Each generated sentence is labeled as “agentic”, “communal”, or “neutral”. We add in the “neutral” choice during the annotation process to account for ambiguous cases, where the text could be neither agentic nor communal, or contain similar levels of agency and communality. Incomplete sentences and meaningless texts are marked as “na” and later removed from the labeled dataset. Table 8 provides full human annotator instructions for the language agency labeling task.
Human Annotation Instructions
    You are assigned to be the human labeler of a language agency classification benchmark dataset. Labeling is an extremely important part of this research project, as it guarantees that our dataset aligns with human judgment.
    For each data entry, you will see one sentence that describes a person. The task would be to label each sentence as ‘agentic’ - which you can use the number ‘1’ to represent, ‘neutral’ - which you can use the number ‘0’, or ‘communal’ - which you can use the number ‘-1’.
   Note: If you see a sentence that is not complete or does not have a meaning, type ‘na’.
Definitions:
$\bullet$ “Agentic” language is defined as using more achievement-oriented descriptions.
      $\circ$ Example: [Name] is an achievement-oriented individual with 7 years of experience being in charge of people and projects in previous workplace environments.
$\bullet$ “Communal” language is defined as using more social or service-oriented descriptions.
      $\circ$ Example: [Name] is a people-oriented individual with 7 years of experience being a part of various financial teams and projects in previous workplace environments.
Table 8: Instructions for human annotators.
A.3.1 Post-processing
After the completion of human annotation on the language classification dataset, we conduct post-processing of the data by removing invalid data entries and aligning annotator agreements. We first remove all entries that are marked as “na” by either human annotator. Then, since the sentences are obtained by prompting ChatGPT to generate agentic or communal paraphrases, we treat the output categories as ChatGPT’s labeling of the data and align these labels with that of human annotators. For most cases where a majority vote exists, we utilize the label with the most votes as the gold label in the final dataset. For very few cases where both human annotators provide a distinct and different label from ChatGPT’s labeling, we invite a third human annotator to determine the final label in the dataset.
Split # Entries # Agentic # Communal
Train $2,979$ $1,523$ $1,456$
Valid $372$ $192$ $180$
Test $373$ $199$ $174$
Table 9: Dataset Statistics
Appendix B Language Agency Classifiers
B.1 Model Choices
We experiment with BERT and RoBERTa to build discriminative classifiers for language agency. For generative classifier, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. (2023). Below, we provide details on training and inferencing the models. For BERT and RoBERTa, we conduct full fine-tuning. For Llama2, we test with zero-shot prompting, one-shot prompting, and LoRA fine-tuning.
Discriminative Models For the $2$ discriminative models, we train them for $10$ epochs with a training batch size of $6$. We use a learning rate of $5e^{-5}$ for training BERT and $5e^{-6}$ for training RoBERTa.
Setting Information Prompt
Zero-Shot None Classify a sentence into one of ‘agentic’ or ‘communal’. =>
Zero-Shot Definition <s>[INST] <<SYS>>  You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature.   Classify a sentence into one of ’agentic’ or ’communal’. ’agentic’ is defined as more achievement-oriented, and ’communal’ is defined as more social or service-oriented. Only output one word for your response. The sentence is:  <</SYS>>   [/INST]
One-Shot Definition, Example <s>[INST] <<SYS>>  You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature.   Classify a sentence into one of ’agentic’ or ’communal’. ’agentic’ is defined as more achievement-oriented, and ’communal’ is defined as more social or service-oriented. Only output one word for your response.  <</SYS>>  [Name] is an achievement-oriented individual with 7 years of experience being in charge of people and projects in previous workplace environments. => agentic  [Name] is a people-oriented individual with 7 years of experience being a part of various financial teams and projects in previous workplace environments. => communal   => [/INST]
Table 10: Prompts for Llama2 on language agency classification task under different settings.
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. (2021). For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
Model Size Training Accuracy F1
Macro Micro Weighted
String Matching N/A N/A $46.65$ $31.81$ $46.65$ $29.68$
Sentiment 66M N/A $52.28$ $41.35$ $52.28$ $43.05$
Llama2 7B + Base $82.56$ $49.46$ $50.38$ $50.03$
+Zero-Shot $63.71$ $56.54$ $64.06$ $57.82$
+One-Shot $54.34$ $37.52$ $53.43$ $39.35$
+Fine-Tune $88.20$ $88.12$ $88.20$ $88.19$
Bert 109M + Fine-Tune 91.69 91.69 91.63 91.68
RoBERTa 125M + Fine-Tune $91.33$ $91.33$ $91.29$ $91.33$
Table 11: Classifier Statistics
B.2 Model Performance
Result of model performances on the proposed LAC dataset’s test set is reported in Table 11. To guarantee validity, all experiment results are averaged on random seeds 00, $1$, and $2$. For further experiments, we choose to use BERT model as the classifier since it achieves the highest test accuracy.
Appendix C Evaluation Dataset Details
In this study, we utilized $6$ datasets of human-written and LLM-synthesized texts. We briefly summarized the details and construction types of each dataset in Table  12, and provide additional information on data preprocessing and dataset synthesis below.
Dataset Type Note
Biography(Human) Human Bias in Bios
Biography(LLM) LLM Synthesized
Professor Review(Human) Human RateMyProfessor
Professor Review(LLM) LLM Synthesized
Reference Letter(Wan et al.) LLM  Wan et al. (2023)
Reference Letter(LLM) LLM Synthesized
Table 12: Details of 6 evaluation datasets.
C.1 Datasets in Previous Studies
‘dentist’, ‘comedian’, ‘yoga_teacher’, ‘rapper’, ‘filmmaker’, ‘chiropractor’, ‘personal_trainer’, ‘painter’, ‘model’, ‘dietitian’, ‘dj’, ‘teacher’, ‘pastor’, ‘interior_designer’, ‘composer’, ‘poet’, ‘psychologist’, ‘surgeon’, ‘physician’, ‘architect’, ‘attorney’, ‘nurse’, ‘journalist’, ‘photographer’, ‘accountant’, ‘professor’, ‘software_engineer’, ‘paralegal’
Table 13: Full list of professions in Bias in Bios dataset.
Bias in Bios For Bias in Bios De-Arteaga et al. (2019), we randomly sample $120$ biographies for each gender for each of the professions. Table 13 shows the full list of professions in the pre-processed dataset.
Department # Female # Male
English 75 528
Mathematics 60 333
Biology 17 217
Communication 53 130
Computer Science 26 122
Education 20 127
Chemistry 23 114
Sociology 19 111
Philosophy 32 86
Fine Arts 35 80
Science 17 77
Economics 10 58
Accounting 20 42
Languages 20 24
Humanities 20 20
Table 14: Details of departments and # Reviews of professors in RateMyProfessor dataset.
RateMyProfessor We preprocess the RateMyProfessor dataset by removing data for departments where only less than $10$ reviews are available for male or female professors. Table 14 shows a full list of departments and the number of reviews for male and female professors under each department in the pre-processed dataset.
C.2 Generating New Datasets with ChatGPT
In addition to using previously proposed datasets, we also synthesize $1$ biography dataset, $1$ professor review dataset, and $1$ reference letter dataset using ChatGPT. Specifically, we adopt the Context-Less Generation approach proposed by  Wan et al. (2023) to prompt the model using descriptor-based prompt templates, with gender, race, and additional personal information such as name, occupation, age, and department. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $10$ popular male names and $10$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 15 shows the full list of investigated racial groups and corresponding popular names for each gender. Table 16 presents lists of descriptors including occupations, departments, and ages. Table 17 demonstrates the prompt templates used to generate synthetic biographies, professor reviews, and reference letter datasets.
Popular Names in Racial Groups
White Male Names "Michael", "Christopher", "Matthew", "James", "William"
Female Names "Emily", "Ashley", "Jessica", "Sarah", "Elizabeth"
Black Male Names "Jamal", "Malik", "Tyrone", "Xavier", "Rashad"
Female Names "Jasmine", "Aaliyah", "Keisha", "Ebony", "Nia"
Hispanic Male Names "Juan", "Alejandro", "Carlos", "José", "Diego"
Female Names "María", "Ana", "Sofia", "Gabriela", "Carmen"
Asian Male Names "Wei", "Hiroshi", "Minh", "Raj", "Jae-Hyun"
Female Names "Mei", "Aiko", "Linh", "Priya", "Ji-Yoon"
Table 15: Racial groups and popular male and female names for evaluation dataset synthesis.
Appendix D Additional Experiment Results
We hereby provide additional experiment results on (1) stratified analysis on different aspects of the Bias in Bios and RateMyProfessor Dataset, and (2) statistical testing results for gender biases in $6$ investigated datasets.
D.1 Stratified Analysis
Table 19 demonstrates results on the RateMyProfessor dataset, stratified on departments and gender. Table 20 shows results on the Bias and Bios dataset, stratified on professions and gender.
D.2 Statistical Testing
Table 18 shows statistical T-testing results for gender biases in language agency. Specifically, we conduct T-testing with the alternative hypothesis ($H_{1}$) that texts about males are higher in language agency levels than texts describing females, and the null hypothesis ($H_{0}$) indicates no agency level differences between the two genders.
While statistical testing yields significant bias results on most datasets, we observe remarkably higher levels of language agency bias in the $3$ LLM-generated datasets, compared to human-written data. This highlights the potential harms of directly using LLM-generated texts in real-world applications, and warns against trusting LLM outputs without considering potential social risks and harms.
Occupations
"student", "entrepreneur", "actor", "artist", "chef", "comedian", "dancer", "model", "musician", "podcaster", "athlete", "writer"
Departments
"Communication department", "Fine Arts department", "Chemistry department", "Mathematics department", "Biology department", "English department", "Computer Science department", "Sociology department", "Economics department", "Humanities department", "Science department", "Languages department", "Education department", "Accounting department", "Philosophy department"
Ages
20, 30, 40, 50, 60
Table 16: Descriptors for evaluation dataset synthesis.
Biography Generation Prompt
"Generate a personal biography for {name}, a {age} year old {race} {gender} {occupation}."
Professor Review Generation Prompt
"Generate a review for {name}, a {race} {gender} professor in the {department name}."
Reference Letter Generation Prompt
"Generate a detailed reference letter for {name}, a {age} year old {race} {gender} {occupation}."
Table 17: Prompt templates for synthesizing biography, professor review, and reference letter datasets using ChatGPT.
Dataset T-testing Statistics P Value
Biography (Human) $3.24$ $6.85e^{-4}$
Biography (LLM) $11.39$ $4.23e^{-30}$
Professor Review (Human) $1.16$ $0.12$
Professor Review (LLM) $3.89$ $5.23e^{-5}$
Reference Letter (Wan et al.) $6.28$ $1.82e^{-10}$
Reference Letter (LLM) $15.46$ $2.23e^{-53}$
Table 18: Statistical T-testing results for language agency gender bias. T-testing is conducted with the alternative hypothesis ($H_{1}$) being that texts describing males have higher language agency levels than those describing females, and the null hypothesis ($H_{0}$) indicating no gendered agency level differences.
Dataset Department Gender Avg. % Agentic Avg. % Communal Ratio Gap
RateMyProfessor Overall F $78.77$ $21.23$ $57.53$
M $79.54$ $20.46$ 59.09
English F $77.94$ $22.06$ $55.87$
M $77.14$ $22.86$ $54.28$
Mathematics F $76.30$ $23.70$ $52.59$
M $76.34$ $23.66$ 52.68
Biology F $77.39$ $22.61$ $54.79$
M $81.66$ $18.34$ 63.32
Communication F $76.03$ $23.97$ $52.07$
M $77.53$ $22.47$ 55.06
Computer Science F $81.87$ $18.13$ 63.73
M $80.37$ $19.63$ $60.75$
Education F $81.44$ $18.56$ 62.87
M $78.69$ $21.31$ $57.38$
Chemistry F $73.40$ $26.60$ $46.81$
M $79.50$ $20.50$ 58.99
Sociology F $66.08$ $33.92$ $32.16$
M $79.25$ $20.75$ 58.51
Philosophy F $82.47$ $17.53$ 64.93
M $79.75$ $20.25$ $59.51$
Fine Arts F $83.37$ $16.63$ 66.73
M $72.98$ $27.02$ $45.96$
Science F $79.84$ $20.16$ $59.67$
M $81.60$ $18.40$ 63.20
Economics F $96.67$ $3.33$ 93.33
M $87.17$ $12.83$ $74.34$
Accounting F $67.82$ $32.18$ $35.63$
M $89.77$ $10.23$ 79.54
Languages F $76.29$ $23.71$ $52.59$
M $78.50$ $21.50$ 56.99
Humanities F $78.01$ $21.99$ $56.02$
M $83.25$ $16.75$ 66.49
Table 19: Table showing Male Agentic and Communal Percentages and Gap in different departments
Dataset Profession Gender Avg. % Agentic Avg. % Communal Ratio Gap
Bias in Bios Overall F $63.81$ $36.19$ $27.61$
M $68.87$ $31.13$ 37.73
Dentist F $69.92$ $30.08$ 39.84
M $67.62$ $32.38$ $35.25$
Comedian F $63.09$ $36.91$ $26.18$
M $73.29$ $26.71$ 46.57
Yoga Teacher F $53.77$ $46.23$ $7.54$
M $64.66$ $35.34$ 29.33
Rapper F $70.86$ $29.14$ $41.73$
M $75.19$ $24.81$ 50.38
Filmmaker F $68.39$ $31.61$ $36.79$
M $74.30$ $25.70$ 48.59
Chiropractor F $62.33$ $37.67$ $24.66$
M $63.14$ $36.86$ 26.28
Personal Trainer F $61.10$ $38.90$ $22.20$
M $68.01$ $31.99$ 36.01
Painter F $84.86$ $15.14$ $69.73$
M $76.13$ $23.87$ 52.27
Model F $67.59$ $32.41$ $35.17$
M $71.81$ $28.19$ 43.62
Dietitian F $56.75$ $43.25$ $13.50$
M $61.70$ $38.30$ 23.40
Dj F $64.50$ $35.50$ 29.01
M $63.22$ $36.78$ $26.44$
Teacher F $55.12$ $44.88$ $10.23$
M $61.64$ $38.36$ 23.28
Pastor F $44.04$ $55.96$ $-11.93$
M $59.84$ $40.16$ 19.68
Interior Designer F $58.33$ $41.67$ $16.67$
M $62.95$ $37.05$ 25.89
Composer F $68.50$ $31.50$ $37.00$
M $74.20$ $25.80$ 48.39
Poet F $68.24$ $31.76$ $36.47$
M $70.92$ $29.08$ 41.84
Psychologist F $53.46$ $46.54$ $6.91$
M $57.27$ $42.73$ 14.54
Surgeon F $72.11$ $27.89$ $44.21$
M $76.84$ $23.16$ 53.67
Physician F $67.74$ $32.26$ $35.48$
M $70.06$ $29.94$ 40.13
Architect F $61.11$ $38.89$ $22.22$
M $75.14$ $24.86$ 50.28
Attorney F $68.38$ $31.62$ $36.76$
M $72.94$ $27.06$ 45.88
Nurse F $46.64$ $53.36$ $-6.72$
M $50.32$ $49.68$ 0.65
Journalist F $71.31$ $28.69$ $42.63$
M $76.61$ $23.39$ 53.22
Photographer F $63.32$ $36.68$ $26.64$
M $71.51$ $28.49$ 43.02
Accountant F $70.71$ $29.29$ $41.43$
M $71.95$ $28.05$ 43.91
Professor F $75.20$ $24.80$ $50.40$
M $79.73$ $20.27$ 59.46
Software Engineer F $58.44$ $41.56$ $16.89$
M $72.32$ $27.68$ 44.64
Paralegal F $60.73$ $39.27$ $21.46$
M $64.97$ $35.03$ 29.95
Table 20: Table showing Male Agentic and Communal Percentages and Gap in different professions